OpenStudy (anonymous):

Find the curve the length of the curve define by r= <2t^3/2, cos 2t, sin 2t >, 0=

OpenStudy (anonymous):

differentiate the r with respect to t let me know what you get . r'(t) = ???

OpenStudy (anonymous):

i think is r' = <1/2 t^1/2, -2sin 2t, 2 cos 2t>?

OpenStudy (anonymous):

check it again i think first term shuld be 3t^1/2

OpenStudy (anonymous):

yes is 3 ^1/2 cuz 2 3/2 t^ 3/2 -1

OpenStudy (anonymous):

now magnitude?

OpenStudy (anonymous):

ok good take mangitude of r'(t) .

OpenStudy (anonymous):

sqrt(3 t^1/2)^2+(2sin 2t)^2+ (2 cos 2t)^2 (9 t^1/4)+4sin^2 2t+ 4cos^2 2t

OpenStudy (anonymous):

this part is were im stuck

OpenStudy (anonymous):

so far i know that sin^2 + cos^2=1 but 4sin^2 2t+ 4cos^2 2t

OpenStudy (anonymous):

still cos^2(2t)+sin^2(2t) =1 just take out 4 as common !

OpenStudy (anonymous):

ok i see @sami-21 4 (cos^2(2t)+sin^2(2t))

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

sqrt (9 t^1/4)+4= 3 t^1/2+2

OpenStudy (anonymous):

check it again !!! you cannot take square root like that .

OpenStudy (anonymous):

k

OpenStudy (anonymous):

is 3 t^1/8 + 2?

OpenStudy (anonymous):

\[\Large |r'(t)|= \sqrt{(3t^\frac{1}{2})^2+(2\sin(2t)^2+(2\cos(2t)^2}\] simplify \[\Large |r'(t)|=\sqrt{9t+4}\]

OpenStudy (anonymous):

what happen to the 1/2?

OpenStudy (anonymous):

it is squared !

OpenStudy (anonymous):

one minute 1/2 is the same as sqrt

OpenStudy (anonymous):

yes it is same .

OpenStudy (anonymous):

got it so | r'(t)| =3t+4

OpenStudy (anonymous):

there is sqrt as well :)

OpenStudy (anonymous):

now integrate |r'(t)| fom 0..1

OpenStudy (anonymous):

\[\Large Length=\int\limits_{0}^{1} \sqrt{3t+4}dt\]

OpenStudy (anonymous):

is \[\sqrt{3t+2}?\]

OpenStudy (anonymous):

cuz 4 is perfect squared

OpenStudy (anonymous):

sorry i did mistake .

OpenStudy (anonymous):

because |r'(t)|= sqrt{9t+4} \[\Large Length=\int\limits_{0}^{1}\sqrt{9t+4}dt\] can you integrate this

OpenStudy (anonymous):

chain rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/2 (9t +4)^-1/2 * 9= 9/ 2 sqrt (9t+4)

OpenStudy (anonymous):

now how inegrate this